light of wavelength 589.3nm falls normally on a plane transmission grating having 600 lines/cm. calculate the angle at which the principal maxima of the first order form


1
Expert's answer
2021-02-05T14:43:16-0500

"d\sin\theta=m\lambda"


"d=\frac{0.01}{600}=1.67\cdot10^{-5}\ (m)"


"\theta=\sin^{-1}(m\lambda/d)=\sin^{-1}(1\cdot589.3\cdot10^{-9}/1.67\cdot10^{-5})\approx2°" . Answer